Theory and Applications of Categories, Vol. 27, No. 10, 2012, pp. 210–221.

DESCENT IN MONOIDAL CATEGORIES

BACHUKI MESABLISHVILI

Abstract. We consider a symmetric monoidal closed category 𝒱 = (𝒱 , ⊗, I, [−, −]) together with a regular injective object Q such that the functor [−, Q]: 𝒱 → 𝒱 ^{op} is comonadic and prove that in such a category, as in the monoidal category of abelian groups, a morphism of commutative monoids is an effective descent morphism for modules if and only if it is a pure monomorphism. Examples of this kind of monoidal categories are elementary toposes considered as cartesian closed monoidal categories, the module categories over a commutative ring object in a Grothendieck topos and Barr’s star-autonomous categories.

1. Introduction

Grothendieck’s descent theory for modules in a symmetric monoidal category 𝒱 = (𝒱 , ⊗, I) is the study of which morphisms ι : A → B of commutative 𝒱 -monoids are effective descent morphisms in the sense that the corresponding extension-of-scalars functor B ⊗_{A} − : _{A}𝒱 → _{B}𝒱 from the category of (left) A-modules to the category of (left) B-modules is comonadic. In [10], [11] and [12], we looked at the case where 𝒱 is the monoidal category of abelian groups, or a star-autonomous category in the sense of Barr [1] and proved that a morphism ι : A → B of commutative 𝒱 -monoids is an effective descent morphism for modules if and only if it is a pure morphism in _{A}𝒱 (that is, for any A-module V, the morphism

ι ⊗_{A} V : V = A ⊗_{A} V → B ⊗_{A} V

is a regular monomorphism). The aim of this paper is to provide a unifying categorical approach to these results. Explicitly the setting in which we work is a symmetric monoidal closed category 𝒱 = (𝒱 , ⊗, I, [−, −]) together with a regular injective object Q such that the functor [−, Q]: 𝒱 → 𝒱 ^{op} is comonadic. Our approach is based on the observation that the proof given in [10] makes heavy use of the description of purity by means of the functor Hom_{ℤ}(−, ℚ/ℤ) : Ab → Ab^{op} which is conservative and preserves all coequalizers, and thus is, in particular, comonadic. In the case of Barr’s star-autonomous categories, the corresponding functor is an equivalence of categories.

The work was partially supported by Volkswagen Foundation (Ref.: I/85989) and Shota Rustaveli National Science Foundation Grant DI/12/ 5-103/11.

Received by the editors 2012-03-26 and, in revised form, 2012-10-03.

Published on 2012-10-09 in the volume of articles from CT2011.

2010 Mathematics Subject Classification: 18A20, 18D10, 18D35.

Key words and phrases: symmetric monoidal categories, effective descent morphisms, pure morphisms.

\begin{matrix}c \\ ◯ \end{matrix} Bachuki Mesablishvili, 2012. Permission to copy for private use granted.

In the first section, we recall some elementary facts about modules in a symmetric monoidal closed category. In section 2, we prove our main result, and in Section 3, we apply it to elementary toposes, the module categories over a commutative ring object in a Grothendieck topos and Barr’s star-autonomous categories.

As background to the subject, we refer to S. Mac Lane [7] for generalities on category theory and to G. Janelidze and W. Tholen [3], [4] and [5] for descent theory.

2. Preliminaries

We begin by recalling from [7] and [13] some elementary facts about modules in a symmetric monoidal closed category.

Suppose that 𝒱 is a fixed symmetric monoidal closed category with tensor product ⊗, unit object I, and internal-hom [−, −]; recall that 𝒱 is closed means that each functor V ⊗ − : 𝒱 → 𝒱 has a right adjoint [V, −] : 𝒱 → 𝒱 . Recall further that the adjunction V ⊗ − ⊣ [V, −] is internal, in the sense that one has natural isomorphisms

[V ⊗ W, Y] ≃ [W, [V, Y]]  (1)

For simplicity of exposition we treat ⊗ as strictly associative and I as a strict unit, which is justified by Mac Lane’s coherence theorem [7] asserting that every monoidal category is equivalent to a strict one.

A monoid in 𝒱 (or 𝒱 -monoid) consists of an object A of 𝒱 endowed with a multiplication m_{A} : A ⊗ A → A and unit morphism e_{A} : I → A such that the usual identity and associative conditions are satisfied. A monoid is called commutative if the multiplication map is unchanged when composed with the symmetry. We write Mon(𝒱 ) for the category of 𝒱 -monoids.

Recall further that, for any 𝒱 -monoid A = (A, e_{A}, m_{A}), a left A-module is a pair (V, ρ_{V}), where V is an object of 𝒱 and ρ_{V} : A⊗V → V is a morphism in 𝒱 , called the action (or the A-action) on V, such that ρ_{V} (m_{A} ⊗ V) = ρ_{V} (A⊗ρ_{V}) and ρ_{V} (e_{A}⊗V) = 1.

For a given 𝒱 -monoid A, the left A-modules are the objects of a category _{A}𝒱 . A morphism f : (V, ρ_{V}) → (W, ρ_{W}) is a morphism f : V → W in 𝒱 such that ρ_{W} (A⊗f) = fρ_{V}. Analogously, one has the category 𝒱_{A} of right A-modules.

The forgetful functor _{A}U: _{A}𝒱 → 𝒱 that takes a left A-module (V, ρ_{V}) to the object V has a left adjoint _{A}F : 𝒱 → _{A}𝒱 sending an object V ∈ 𝒱 to the ”free” A-module (A⊗V, A⊗ρ_{V}).

There is another way of representing the category of left A-modules that involves algebras over the monad associated to the 𝒱 -monoid A.

Every 𝒱 -monoid A = (A, e_{A}, m_{A}) defines a monad L(A) = (T, η, μ) on 𝒱 by

• T(V) = A ⊗ V,

• η_{V} = e_{A} ⊗ V : V → A ⊗ V,

• μ_{V} = m_{A} ⊗ V : A ⊗ A ⊗ V → A ⊗ V.

It is well known that the corresponding Eilenberg-Moore category 𝒱 ^{L(A)} of L(A)-algebras is exactly the category _{A}𝒱 of left A-modules, and that _{A}U ⊣ _{A}F is the familiar forgetfulfree adjunction between 𝒱 ^{L(A)} and 𝒱 . This gives in particular that the forgetful functor AU: _{A}𝒱 → 𝒱 is monadic. Hence the functor _{A}U creates those limits that exist in 𝒱 . Moreover, since the functor A⊗− : 𝒱 → 𝒱 admits as a right adjoint the functor [A, −]− : 𝒱 → 𝒱 , the forgetful functor _{A}U has a right adjoint sending an object V ∈ 𝒱 to the object [A, V], where [A, V] is an object of _{A}𝒱 via the transpose A⊗[A, V] → [A, V] of

the composite A⊗A⊗[A, V] ⟶^{m_{A}⊗[A,V]} A⊗[A, V] ⟶^{ev_{V}} V, where ev_{V} : A⊗[A, V] → V is

the V -component of the counit of the adjunction A⊗− ⊣ [A, −]. In particular, _{A}U creates those colimits that exist in 𝒱 .

If 𝒱 admits coequalizers, A is a 𝒱 -monoid, (V, ϱ_{V}) ∈ 𝒱_{A} a right A-module, and (W, ρ_{W}) ∈ _{A}𝒱 a left A-module, then their tensor product (over A) is the object part of the following coequalizer

\xymatrix{V ⊗A ⊗W \ar[r]^{ϱ_{V} ⊗W} \ar[r]_{V ⊗ρ_{W}} & V ⊗W \ar[r] & V ⊗_{A}W.}

When A is commutative, then for any (V, ρ_{V}) ∈ _{A}𝒱 , the composite ρ′_{V} = ρ_{V} τ_{V,A} : V ⊗A → V, where τ is the symmetry for 𝒱 , defines a right A-action on V. Similarly, if (W, ϱ_{V}) ∈ 𝒱_{A}, then ϱ′_{W} = ϱ_{V} τ_{W,A} : W ⊗A → W defines a left A-action on W. These two constructions establish an equivalence between _{A}𝒱 and 𝒱_{A}, and thus we do not have to distinguish between left and right A-modules. In this case, the tensor product of two A-modules is another A-module, and tensoring over A makes _{A}𝒱 (as well as 𝒱_{A}) into a symmetric monoidal category with unit A. If, in addition, 𝒱 admits equalizers, then this monoidal structure on _{A}𝒱 is closed: The internal Hom-object _{A}[V, W] of two A-modules defined to be the equalizer in 𝒱 of

\xymatrix{[V, W] \ar[r] \ar[r] & [A ⊗ V, W],}

where one of the morphisms is induced by the action of A on V, and the other is the composition of A ⊗ − : [V, W] → [A ⊗ V, A ⊗ W] followed by the morphism induced by the action of A on W.

In what follows, 𝒱 denotes a fixed symmetric monoidal closed category with equalizers and coequalizers.

3. Descent theory in monoidal categories

3.1. Let us recall that a morphism in a category A is a regular monomorphism if it is an equalizer of some pair of morphisms. Recall also that a regular injective object in A is an object X ∈ A which has the extension property with respect to regular monomorphisms; that is, if every extension problem

\xymatrix{A \ar[d]_{f} \ar@{^{(}->}[r]^{m} & B \ar@{.>}[dl]^{f̄} \\ X}

with m a regular monomorphism has a solution f̄ : B → X extending f along m, i.e., satisfying f̄m = f.

3.2. A pointed 𝒱 –endofunctor on 𝒱 is a pair (T, η), where T : 𝒱 → 𝒱 is a 𝒱 – endofunctor on 𝒱 and η : 1 → T is a 𝒱 –natural transformation. Let (T, η) be a pointed 𝒱 – endofunctor on 𝒱 . For an object Q of 𝒱 , we get from T a functor

[T (−), Q]: 𝒱 → 𝒱 ^{op},

and we can consider the natural transformation

[η_{−}, Q]: [T (−), Q] → [−, Q].

3.3. Proposition [12] The natural transformation [η_{−}, Q] is a split epimorphism if and only if the morphism η_{Q}: Q → T (Q) is a split monomorphism. In particular, if Q is a regular injective object in 𝒱 , then the natural transformation [η_{−}, Q] is a split epimorphism if and only if η_{Q} is a regular monomorphism in 𝒱 .

Recall [8] that a monad T on a category A is called of descent type if the free T-algebra functor F^{T} : A → A^{T} is precomonadic, and T is called of effective descent type if F^{T} is comonadic.

3.4. Theorem Let 𝒱 have a regular injective object Q such that the functor

[−, Q]: 𝒱 → 𝒱 ^{op}

is comonadic. For any commutative monoid A = (A, e_{A}, m_{A}) in 𝒱 , the following are equivalent:

(i) the morphism e_{A}: I → A is pure; that is, for any object V ∈ 𝒱 , the morphism

e_{A} ⊗ V : V → A ⊗ V

is a regular monomorphism;

(ii) the morphism e_{A}⊗ Q : Q → A ⊗ Q is a regular monomorphism;

(iii) the natural transformation [e_{A}⊗ −, Q] is a split epimorphism;

(iv) the morphism [e_{A}, Q] : [A, Q] → [I, Q] is a split epimorphism;

(v) the monad L(A) is of descent type;

(vi) the monad L(A) is of effective descent type.

Proof. Clearly, (i) always implies (ii), while (ii) and (iii) are equivalent by Proposition 3.3, since the pair (A ⊗ −, e_{A}⊗ −) is a pointed 𝒱 –endofunctor on 𝒱 .

Since the I-component of the natural transformation [e_{A}⊗ −, Q] is just the morphism [e_{A}, Q] : [A, Q] → [I, Q], (iii) implies (iv).

To see that (iv) implies (vi), note first that to say that the natural transformation [e_{A}⊗ −, Q] is a split epimorphism is to say that the monad L(A) is [−, Q]–separable [8]. Next, observe that the diagram

\xymatrix{𝒱 \ar[r]^{A⊗−} \ar[d]_{[−,Q]} & 𝒱 \ar[d]^{[−,Q]} \\ 𝒱 ^{op} \ar[r]_{[A,−]^{op}} & 𝒱 ^{op}}

commutes up to natural isomorphism by (1). Now, since 𝒱 ^{op} admits equalizers (and hence is Cauchy complete) and since the functor 𝒱 (−, Q): 𝒱 → 𝒱 ^{op} is comonadic, one can apply [8, Theorem 3.22] to the diagram to conclude that the monad L(A) is of effective descent type.

(vi) trivially implies (v), while the implication (v) ⇒ (i) follows from [8, Theorem 2.3 (i)].

3.5. Lemma Let 𝒱 have an object Q such that the functor

𝒱 (−, Q): 𝒱 → 𝒱 ^{op}

is comonadic, and let A = (A, e_{A}, m_{A}) be a commutative monoid in 𝒱 . Write Q_{A} for the object [A, Q] of 𝒱 . Then Q_{A} ∈ _{A}𝒱 and the functor

A[−, Q_{A}]: _{A}𝒱 → (_{A}𝒱 )^{op}

is comonadic. Moreover, if Q is regular injective in 𝒱 , then Q_{A} is regular injective in A𝒱 .

Proof. We have already seen that the functor _{A}U: _{A}𝒱 → 𝒱 is left adjoint, with right adjoint [A, −]: 𝒱 → _{A}𝒱 . Hence, every [A, V] (in particular, [A, Q]) is an object of the category _{A}𝒱 . Thus Q_{A} ∈ _{A}𝒱 .

Since for any V ∈ _{A}𝒱 , _{A}[V, [A, Q]] ≃ [A ⊗_{A} V, Q] ≃ [V, Q] (see, for instance, [13]), one has commutativity (up to isomorphism) in

\xymatrix{_{A}𝒱 \ar[r]^{_{A}[−,Q_{A}]} \ar[d]_{_{A}U} & (_{A}𝒱 )^{op} \ar[d]^{(_{A}U)^{op}} \\ 𝒱 \ar[r]^{[−,Q]} & 𝒱 ^{op}}

and since • the functor _{A}[−, Q_{A}] : _{A}𝒱 → (_{A}𝒱 )^{op} admits as a right adjoint the functor _{A}[−, Q_{A}] :

(_{A}𝒱 )^{op} → _{A}𝒱 ,

• the functor _{A}U: _{A}𝒱 → 𝒱 preserves all small limits, and thus, in particular, equalizers of _{A}[−, Q_{A}]–split pairs;

• the functor [−, Q]: 𝒱 → 𝒱 ^{op}, being comonadic, preserves equalizers of [−, Q]–split pairs;

• the functor _{A}U (and hence also (_{A}U)^{op}) is conservative,

it follows from the dual of [9, Theorem 5.5] that the functor

A[−, Q_{A}]: _{A}𝒱 → (_{A}𝒱 )^{op}

is comonadic.

Now, using that

• the functor [A, −]: 𝒱 → _{A}𝒱 is right adjoint to the functor _{A}U: _{A}𝒱 → 𝒱 ;

• _{A}U preserves regular monomorphisms;

• Q is regular injective in 𝒱 ,

it is easy to show that the object Q_{A} = [A, Q] is regular injective in _{A}𝒱 . This completes the proof.

3.6. For any symmetric monoidal closed category 𝒱 , we denote the category of commutative monoids in 𝒱 by CMon(𝒱 ). It is well-known that for any commutative 𝒱 -monoid A, the co-slice category A/CMon(𝒱 ) is equivalent to the category CMon(_{A}𝒱 ). In other words, to give a commutative monoid B in the symmetric monoidal closed category _{A}𝒱 is to give a morphism A → B of commutative monoids in 𝒱 . The latter morphism serves as the unit morphism of the _{A}𝒱 -monoid B. If ι: A → B is a morphism in CMon(𝒱 ), then the corresponding commutative monoid in 𝒱_{A} will be denoted by B_{ι}.

One says that a morphism ι : A → B of commutative 𝒱 -monoids is an (effective) descent morphism if the functor B⊗_{A}− : _{A}𝒱 → _{B}𝒱 is precomonadic (comonadic).

Identifying the morphism ι : A → B with the monoid B_{ι} in the monoidal category A𝒱 and considering the monad L(B_{ι}) = (T_{ι}, η_{ι}, μ_{ι}) on _{A}𝒱 induced by B_{ι} (thus, T_{ι} = B ⊗_{A} −, η_{ι} = ι ⊗_{A} − and μ_{ι} = m′_{B} ⊗_{A} −, where m′_{B} : B ⊗_{A} B → B is the unique morphism through which m_{B} factors), the category _{B}𝒱 can be seen as the Eilenberg-Moore category of L(B_{ι})-algebras. Hence the category _{B_{ι}}(_{A}𝒱 ) can be identified with the category _{B}𝒱 . Modulo this identification, the functor B⊗_{A}− : _{A}𝒱 → _{B}𝒱 corresponds to the functor B_{ι}⊗_{A}− : _{A}𝒱 → _{B_{ι}}(_{A}𝒱 ). Thus the problem of effectiveness of ι is equivalent to the one of the monad L(B_{ι}). Using this, and the fact that there is a natural isomorphism A[−, Q_{A}] ≃ [−, Q], we get from Lemma 3.5 and Theorem 3.4:

3.7. Theorem Let 𝒱 have a regular injective object Q such that the functor

[−, Q]: 𝒱 → 𝒱 ^{op}

is comonadic, and let ι: A → B be a morphism of commutative monoids in 𝒱 . The following are equivalent:

(i) ι: A → B is an effective descent morphism;

(ii) ι: A → B is a pure morphism in _{A}𝒱 ; that is, for any A-module V, the morphism

ι ⊗_{A} V : V = A ⊗_{A} V → B ⊗_{A} V

is a regular monomorphism;

(iii) the morphism [ι, Q] : [B, Q] → [A, Q] is a split epimorphism in _{A}𝒱 ;

(iv) the monad L(B_{ι}) is of descent type;

(v) the monad L(B_{ι}) is of effective descent type.

4. Applications

4.1. Monoid modules in an elementary topos Let E be an elementary topos, considered as a cartesian monoidal category. It is well-known [6] that the functor

Ω^{(−)} : E^{op} → E,

where Ω is the subobject classifier for E, is monadic. Hence Ω^{(−)}, seen as a functor E → E^{op}, is comonadic. Moreover, since Ω is an injective object in E (e.g., [6]) and since in E regular monomorphisms coincide with monomorphisms, Theorem 3.7 gives the following result:

4.2. Theorem Let E be an elementary topos. A morphism ι: A → B of commutative monoids in E is an effective descent morphism (or, equivalently, the functor B ⊗_{A} − : AE → _{B}E is comonadic) if and only if ι: A → B is a pure morphism in _{A}E.

4.3. The case of the topos of sets Specialize now to the case where E is the topos of sets, 𝔖𝔢𝔱, so that 𝔖𝔢𝔱-monoids are ordinary monoids, and if A such a monoid, then (left and right) A-modules are more commonly called A-actions. Recall that for any left A–action X, the set 𝔖𝔢𝔱(X, 2), where 2 = {0, 1}, is a right A–action under the definition (f · a)(x) = f(a · x) for all a ∈ A, f ∈ 𝔖𝔢𝔱(X, 2) and x ∈ X (see Section 2). Moreover, for any morphism f : X → Y of left A–actions, the function

𝔖𝔢𝔱(f, 2) : 𝔖𝔢𝔱(Y, 2) → 𝔖𝔢𝔱(X, 2)

is a morphism of right A–actions. It is well-known (e.g., [7]) that, when f is injective, then there exists a map

∃_{f} : 𝔖𝔢𝔱(X, 2) → 𝔖𝔢𝔱(Y, 2)

of sets such that 𝔖𝔢𝔱(f, 2) · ∃_{f} = 1. Recall that, for any map χ : X → 2, the map ∃_{f}(χ) : Y → 2 is defined as follows:

(∃_{f}(χ))(y) = \begin{cases}1, & if there is x ∈ X such that χ(x) = 1 and f(x) = y, \\ 0, & otherwise. \end{cases}

4.4. Proposition Let A be a group. Then, for any injective morphism f : X → Y of left A–actions, the map ∃_{f} : 𝔖𝔢𝔱(X, 2) → 𝔖𝔢𝔱(Y, 2) is a morphism of right A–actions.

Proof. We have to show that ∃_{f}(χ) · a = ∃_{f}(χ · a) for all χ ∈ 𝔖𝔢𝔱(X, 2) and all a ∈ A. If y ∈ Y is an arbitrary element, then (∃_{f}(χ) · a)(y) = (∃_{f}(χ))(a · y), and we have

(∃_{f}(χ) · a)(y) = \begin{cases}1, & if there is x ∈ X such that χ(x) = 1 \\ & and f(x) = a · y, \\ 0, & otherwise. \end{cases}

which, since A is a group and since f is a morphism of left A–actions, may be written as

(∃_{f}(χ) · a)(y) = \begin{cases}1, & if there is x ∈ X such that χ(x) = 1 \\ & and f(a^{−1} · x) = y, \\ 0, & otherwise. \end{cases}

On the other hand, we have

(∃_{f}(χ · a))(y) = \begin{cases}1, & if there is x ∈ X such that (χ · a)(x) = 1 \\ & and f(x) = y, \\ 0, & otherwise. \end{cases}

and hence

(∃_{f}(χ · a))(y) = \begin{cases}1, & if there is x ∈ X such that χ(a · x) = 1 \\ & and f(x) = y, \\ 0, & otherwise. \end{cases}

Comparing (∃_{f}(χ) · a)(y) with (∃_{f}(χ · a))(y), we find that they are equal. So ∃_{f}(χ) · a = ∃_{f}(χ · a), and hence f is a morphism of right A–actions.

Since 𝔖𝔢𝔱(f, 2) · ∃_{f} = 1, a corollary follows immediately:

4.5. Corollary Let A be a group and f : X → Y an injective morphism of left A– actions. Then the map

𝔖𝔢𝔱(f, 2) : 𝔖𝔢𝔱(X, 2) → 𝔖𝔢𝔱(Y, 2)

is a split epimorphism of right A–actions.

We are now ready to state and prove the following

4.6. Theorem Let ι : A → B be a morphism of ordinary commutative monoids. If A is an (abelian) group, then ι is an effective descent morphism if and only if it is an injective map.

Proof. One direction is immediate from Theorem 3.7. Conversely, if ι is injective, then the map

𝔖𝔢𝔱(ι, 2) : 𝔖𝔢𝔱(B, 2) → 𝔖𝔢𝔱(A, 2)

is a split epimorphism of right A–actions by Corollary 4.5. But according to Theorem 3.7, ι is an effective descent morphism if and only if the map 𝔖𝔢𝔱(ι, 2) : 𝔖𝔢𝔱(B, 2) → 𝔖𝔢𝔱(A, 2) is a split epimorphism in _{A}𝔖𝔢𝔱, or equivalently (by the commutativity of A), in 𝔖𝔢𝔱_{A}. Thus, ι is an effective descent morphism.

4.7. The category of internal modules over a Grothendieck topos In order to proceed we need the following easy consequence of a variation of Duskin’s theorem (see [2, Theorem 1.3 of Section 9.1]).

4.8. Theorem A left adjoint additive functor F : 𝒜 → ℬ between abelian categories is comonadic if and only if F is conservative and F preserves those monomorphisms whose cokernel-pairs are F -split.

A monoidal category (𝒱 , ⊗, I) is called abelian monoidal if 𝒱 is abelian and the tensor product is an additive bifunctor. An object V of such a category 𝒱 is said to be flat if the functor V ⊗ − : 𝒱 → 𝒱 preserves monomorphisms.

4.9. Proposition Let 𝒱 = (𝒱 , ⊗, I, τ, [−, −]) be a symmetric monoidal closed abelian category. Suppose 𝒱 has a generating family formed by flat objects. If Q ∈ 𝒱 is an injective cogenerator, then the functor

[−, Q] : 𝒱 → 𝒱 ^{op}

is comonadic.

Proof. We first observe that the functor [−, Q] : 𝒱 → 𝒱 ^{op} admits as a right adjoint the functor [−, Q] : 𝒱 ^{op} → 𝒱 .

Next, if f is a morphism in 𝒱 such that the morphism [f, Q] is an isomorphism, then the map 𝒱 (I, [f, Q]) is bijective. Because of the following chain of bijections 𝒱 (I, [f, Q]) ≃ 𝒱 (I ⊗ f, Q) ≃ 𝒱 (f, Q), it follows that the map 𝒱 (f, Q) is also bijective. But since Q is an injective cogenerator, the functor 𝒱 (−, Q) : 𝒱 ^{op} → Set is faithful; thus, it reflects epimorphisms and monomorphisms, and hence isomorphisms. Therefore, f is also an isomorphism. Consequently, the functor [−, Q] is conservative.

We next show that functor [−, Q] preserves monomorphisms. Indeed, if 𝒱 has a generating family {G_{α}} such that each G_{α} is flat and if f : V → V′ is a monomorphism in 𝒱 , then so also is each G_{α} ⊗ f. Then, since Q is injective in 𝒱 , the map

𝒱 (G_{α} ⊗ f, Q) : 𝒱 (G_{α} ⊗ V′, Q) → 𝒱 (G_{α} ⊗ V, Q),

and hence also the map

𝒱 (G_{α}, [f, Q]) : 𝒱 (G_{α}, [V′, Q]) → 𝒱 (G_{α}, [V, Q]),

is surjective for all α. But {G_{α}} is a generating family for 𝒱 , i.e. the family of functors {𝒱 (G_{α}, −) : 𝒱 → Set} is collectively faithful; in particular, this family collectively reflects epimorphisms. Therefore, the morphisms [f, Q] is an epimorphism in 𝒱 . Applying now the dual of Theorem 4.8 gives that the functor [−, Q] is comonadic.

We now consider the symmetric monoidal closed category Ab(E) of internal abelian groups in a Grothendieck topos E. It is well-known that (commutative) monoids in Ab(E) are internal (commutative) rings in E, and that _{A}Ab(E), A ∈ Mon(Ab(E)), is the category Mod_{A}(E) of internal left A-modules in E. Since Ab(E) is an Ab5 category with generators and sufficiently many injective objects (e.g., [6]), it also has an injective cogenerator, say, Q (see, for example, [14, Lemma 7.12]). Now, since free abelian groups in E are flat in Ab(E) and since Ab(E) has a generator that is a free abelian group (see, [6]), one can combine Proposition 4.9 with Theorem 3.7 to conclude the following generalization of the main result of [10] (see also [11]):

4.10. Theorem A morphism ι: A → B of internal commutative rings in a Grothendieck topos E is an effective descent morphism (or, equivalently, the functor B⊗_{A}− : Mod_{A}(E) → Mod_{A}(E) is comonadic) if and only if ι: A → B is a pure morphism of internal (left) A-modules.

4.11. ⋆-autonomous categories Let now 𝒱 be a ∗-autonomous category in the sense of Barr [1]. Then 𝒱 is a symmetric monoidal closed category together with a so-called dualizing object Q such that the adjunction

[−, Q] ⊣ [−, Q]: 𝒱 ^{op} → 𝒱

is an adjoint equivalence. Quite obviously, the functor [−, Q]: 𝒱 → 𝒱 ^{op} is then comonadic. Moreover, it is proved in [12] that any dualizing object is regular injective in 𝒱 if and only if the tensor unit I is regular projective in 𝒱 . Thus, when the tensor unit is regular projective in a ∗-autonomous category, Theorem 3.7 applies (see [12], for more on the descent morphisms in ∗-autonomous categories).

References

[1] M. Barr, ∗-Autonomous Categories, Lecture Notes in Mathematics 752, Springer-Verlag, Berlin (1979).

[2] M. Barr and C. Wells, Toposes, Triples, and Theories, Grundlehren der Math. Wissenschaften 278, Springer-Verlag, 1985.

[3] G. Janelidze and W. Tholen, Facets of Descent, I, Appl. Categorical Structures 2 (1994), 245–281.

[4] G. Janelidze and W. Tholen, Facets of Descent, II, Appl. Categorical Structures 5 (1997), 229–248.

[5] G. Janelidze and W. Tholen, Facets of Descent, III : Monadic Descent for Rings and Algebras, Appl. Categorical Structures 12 (2004), 461–477.

[6] P. Johnstone, Topos Theory, L.M.S. Mathematical Monographs 10, Academic Press, London (1977).

[7] S. Mac Lane, Categories for the Working Mathematician, Graduate Texts in Mathematics Vol. 5, Springer, Berlin-New York (1971).

[8] B. Mesablishvili, Monads of effective descent type and comonadicity, Theory and Applications of Categories 16 (2006), 1–45.

[9] B. Mesablishvili, Descent theory for schemes, Applied Categorical Structures 12 (2004), 485–512.

[10] B. Mesablishvili, Pure morphisms of commutative rings are effective descent morphisms for modules − a new proof, Theory and Applications of Categories 7 (2000), 38–42.

[11] B. Mesablishvili, Pure morphisms are effective for modules, Applied Categorical Structures (in press).

[12] B. Mesablishvili, Descent in ⋆-autonomous categories, Journal of Pure and Applied Algebra 213 (2009), 60–70.

[13] B. Pareigis, Non-additive ring and module theory I. General theory of monoids, Publ. Math. Debrecen 24 (1977), 189–204.

[14] N. Popescu, Abelian Categories with Applications to Rings and Modules, L.M.S. Mathematical Monographs 3, Academic Press, London (1973).

A. Razmadze Mathematical Institute of I. Javakhishvili Tbilisi State University, 2, University St., Tbilisi 0186, and Tbilisi Centre for Mathematical Sciences, Chavchavadze Ave. 75, 3/35, Tbilisi 0168, Republic of Georgia.

Email: bachi@rmi.ge

This article may be accessed at http://www.tac.mta.ca/tac/ or by anonymous ftp at

ftp://ftp.tac.mta.ca/pub/tac/html/volumes/27/10/27-10.{dvi,ps,pdf}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors The typesetting language of the journal is TEX, and L^{A}TEX2e strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors

Clemens Berger, Université de Nice-Sophia Antipolis, cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com

Valeria de Paiva: valeria.depaiva@gmail.com

Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne : kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Glasgow, Tom.Leinster@glasgow.ac.uk Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Edinburgh: Alex.Simpson@ed.ac.uk James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca